OpenStudy (anonymous):

Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid one hundred thirty-five dollars. This year, he paid one hundred sixty-two dollars. What was the percent increase in the cost of his plane ticket? Show your work.

OpenStudy (yanasidlinskiy):

162-135=27 27/135=.2 .2*100=20 20%

OpenStudy (anonymous):

thanks

OpenStudy (yanasidlinskiy):

Your welcome!!:) Anytime!!:)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Isabelle made fifty-six thousand dollars last year. Randall made forty thousand dollars plus fifteen percent commission on one hundred twenty thousand dollars in sales. Who had a higher income last year? Explain.